can someone please help me wirh this wuestion? thanks

Calculate the total number of ATPS from one molecles of fat with 100 carbons with fatty acid of the molecules.

note: fatty acid and glycerol
Answer #1

Since nothing is given in the question about any the glyceride.
So assume that it is a monoacylglycerol which has one glycerol molecule and one fatty acid of 100 carbon.
The fatty acid is degraded into acetyl Co-A inside the mitochondrial matrix.
It required two ATP to transfer fatty acid to the mitochondrial matrix from the cytoplasm.
Fatty acid is degraded into Acrtly Co-A via Beta oxidation.
In each turn of beta-oxidation, two carbon is removed and converted into Acetyl Co-A. During this process, one NADH and one FADH is produced.
To degrade 100 carbon fatty acid, it takes 49 beta-oxidation cycle, which produces 50 acetyl Co-A.
So from 49 cycles -49 NADH and 49 FADH are produced.
Glycerol - Glycerol is converted into dihydroxyacetone phosphate and enters in the glycolysis. It produces one ATP and one NADH to enter glycolysis.
One molecule of Glyceraldehyde 3 phosphate produces 1+1+3 NADH, 1 FADH 3 ATP.
So overall will be 6 NADH + 1 FADH + 4 ATP.
6*2.5 + 1.5+4 = 20.5 ATP.
One acetyl Co-A upon complete oxidation through TCA cycle produces 3 NADH, 1 FADH, and one ATP.
So from 50 acetyl Co-A 3*50 NADH, 50 FADH, and 50 ATP are produced.
So overall NADH = 150+49 = 199
FADH = 50+49 = 99
ATP = 50
One NADH gives 2.5 ATP and one FADH gives 1.5 ATP.
So total ATP = 199*2.5 + 99*1.5 + 50 = 497.5 + 148.5 + 50 = 696 ATP
Glycerol - Glycerol is converted into dihydroxyacetone phosphate and enters in the glycolysis. It produces one ATP and one NADH to enter glycolysis.
One molecule of Glyceraldehyde 3 phosphate produces 1+1+1+3 NADH, 1 FADH, 4 ATP.
So overall will be 6 NADH + 1 FADH + 4 ATP.
6*2.5 + 1.5+4 = 20.5 ATP.
So overall ATP = 20.5+696 = 716.5 ATP.
Two ATP is utilized to transport fatty acid to mitochondria.
So, overall ATP = 716.5 - 2 = 714.5 ATP.
Please write to me if you need more help or clarifications.
Please hit the thumbs up button